OpenStudy (anonymous):

write the given equation in exponential form. log 3 1/9 = -2? show work

OpenStudy (anonymous):

\[\log_{a}x=y,a ^{y}=x\]

OpenStudy (anonymous):

show work @surjithayer

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

\[\log_{3} \frac{ 1 }{ 9 }=x (say)\] \[3^x=\frac{ 1 }{ 9 }=\frac{ 1 }{ 3^2 }=1*3^{-2}\] \[3^x=3^{-2},x=-2\]

OpenStudy (anonymous):

so whats the answer? @surjithayer

OpenStudy (anonymous):

Answer is \[3^{-2}=\frac{ 1 }{ 9 }\]